TWO SERIES OF SQUARE 1-DESIGNS DERIVED FROM CLASSICAL

DESIGNS P G(2, p^{2s}) AND P G_{2}(3, p^{s}) by Vlado Cigić, Mostar, BiH

Abstract.In this paper there are discussed two families of 1-designs derived from classical designs P G(2, p^{2s}) and P G_{2}(3, p^{s}). In the first case we have a series of semisymmetric designs with a divisibility. Mathematics subject classification: 05B25. Keywords: Partial and semi-symmetric design divisibility, resolution, Baer subdesign.